# A computable version of Hall’s Harem Theorem and Geometric von Neumann Conjecture

### v3

@@ 2.2. Reflections
The following notation substantially simplifies the presentation. For any vertex $v\in V$ there exist a vertex from $U$ which is a copy of $v$ (i.e. the same natural number), we denote it by $u_{v}$. If a vertex $u\in U$ has the copy in $V$ then we denote this copy by $v_{u}$.
@@
The graph $\Gamma=(U,V,E)$ is called $U$-reflected if $V$ is a subset of the right copy of $U$ and for every edge $(u,v)\in E$ with $v_{u}\in V$ the edge $(u_{v},v_{u})$ is in $E$ too. If additionally $V$ is a right copy of $U$ then $\Gamma$ is called a fully reflected bipartite graph.
The Theorem 2.5 below states that in the case of fully reflected bipartite graphs the Hall’s harem condition allows us to force some additional properties at the expense of obtaining a perfect $(1,(d-1))$-matching instead of a $(1,d)$-matching. This is an intermediate statement that avoids computability issues as much as it is possible, the proof of this theorem is available in [Dud21]. We will now give necessary details.
@@
Let $f$ be a function. If for some $i\neq 0$ we have $f^{i}(u)=u$ then we will say that $u$ is a periodic point. For such $u$ and the smallest $i\neq 0$ with $f^{i}(u)=u$, we say that $\{u,f(u),\ldots,f^{i-1}(u)\}$ is a cycle of $f$. Any $(1,(d-1))$ matching can be considered as a $(d-1)$ to $1$ function $f:\mathbb{N}\rightarrow\mathbb{N}$. Moreover, if this matching is perfect, such a function $f$ is total and surjective. We roughly want to show that given a fully reflected bipartite graph satisfying Hall’s $d$-harem condition, there is a perfect $(1,(d-1))$ matching $f:\mathbb{N}\rightarrow\mathbb{N}$, such that for each $u$ there exist $i\geq 0$ such that $f^{i}(u)$ is a periodic point.
@@
Let $f:\mathbb{N}\rightarrow\mathbb{N}$ be a $(d-1)$ to $1$ function. We say that $f$ has controlled sizes of its cycles if each of the following conditions holds:
@@
The following theorem is proved in [Dud21].
@@
Then there exist a perfect $(1,d-1)$-matching of $\Gamma$, which realizes a $(d-1)$ to $1$ function $f:\mathbb{N}\rightarrow\mathbb{N}$ with controlled sizes of its cycles.
@@
Before we state a computable version of Theorem 2.5 we describe our approach to computability of classical Hall’s $d$-harem theorem from [DI22a].
@@
To simplify the matter, in the case of bipartite graphs, from now on we will always identify $U$ and $V$ with computable subsets of $\mathbb{N}$ so that $\nu$ is the identity map. Further, admitting that $U\cap V\not=\emptyset$ we distinguish these sets saying that $U$ is taken from the left copy of $\mathbb{N}$ but $V$ is taken from the right one. In particular we often consider graphs where $U=V=\mathbb{N}$.
To formulate a computable version of Hall’s Harem theorem we need the following definition of Kierstead [Kie83].
@@
Manaster and Rosenstein [Man72] showed that there is a computably bipartite, highly computable graph that satisfies Hall’s condition for a bipartite graph to have a $(1,1)$-matching, but has no computable $(1,1)$-matching.
Kierstead [Kie83] found the natural extension of Hall’s condition, called computable expanding Hall’s condition, which allows us to construct computable $(1,1)$-matchings.
Following [DI22a], in this paper we focus on computable perfect $(1,k)$-matchings.
@@
In [DI22a] we introduced the following condition, which implies the existence of computable perfect $(1,k)$-matchings in highly computable bipartite graphs. For $k=1$ it was formulated earlier in [Kie83].
@@
In [DI22a] we have proven the following.
@@ 4.1. Main Theorem and the aim of the construction
In the proof of Theorem A we need a computable perfect $(1,(d-1))$ matching which realizes a function $f:\mathbb{N}\rightarrow\mathbb{N}$, such that the following conditions are satisfied:
@@
    $\{(n,m)\,|\,n\text{ is a periodic point },m\text{ is in the cycle of }n\}$ is computable;
@@
In fact it is easy to see that a computable $(d-1)$ to $1$ function with controlled sizes of its cycles (Definition 2.4) satisfies conditions (I)–(III).
@@
Then there exist a computable perfect $(1,d-1)$-matching of $\Gamma$, which realizes a computable $(d-1)$ to $1$ function $f:\mathbb{N}\rightarrow\mathbb{N}$ with controlled sizes of its cycles.
Main Theorem is proved in the section 8. The next sections focus on a construction used in this proof.
@@ 4.2. Notation used in the construction
  • $M_{n-1}$ is a set of $(1,d-1)$-fans added to $M$ at the end of the $n$-th step. Thus $M=\bigcup\limits_{n=1}^{\infty}M_{n-1}$.
@@
  • $\Gamma^{(n)\perp}$ is a graph with the sets of vertices $(U^{(n)\perp},V^{(n)\perp})$ and the set of edges corresponding to $(1,d-1)$-fans with roots denoted by $u^{\perp}_{i}\in U^{(n)\perp}$.
@@
  • $\Gamma^{(n)\star}:=(U^{(n)}\setminus U^{(n)\perp},V^{(n)}\setminus V^{(n)\perp})$. We will see that $\Gamma^{(n)\star}$ satisfies Hall’s harem condition.
@@
  • For any subgraph $\Gamma^{\prime}$ of $\Gamma^{(n)}$ by $\Gamma^{\prime}(-u^{0}_{n},\ldots,-u^{j}_{n})$ we denote the graph obtained from $\Gamma^{\prime}$ by removal of the $(1,d-1)$-fans of $M_{n}$ with roots $u^{0}_{n},\ldots,u^{j}_{n}$.
  • For any subgraph $\Gamma^{\prime}=(U^{\prime},V^{\prime})$ of $\Gamma^{(n)}$ and any $u_{j}^{\perp}\in U^{(n)\perp}$ by $\Gamma^{\prime}(+u_{j}^{\perp})$ we denote the graph induced in $\Gamma^{(n)}$ by the sets of vertices $U^{\prime}\cup\{u_{j}^{\perp}\}$ and $V^{\prime}\cup\{v^{\perp}_{j,i}:1\leq i\leq d-1\}$.
  • For any subgraph $\Gamma^{\prime}=(U^{\prime},V^{\prime})$ of $\Gamma^{(n)}$ and any vertex $v\in V^{(n)}$ by $\Gamma^{\prime}(+v)$ (resp. $\Gamma^{\prime}(-v)$) we denote the graph induced in $\Gamma^{(n)}$ by the sets of vertices $U^{\prime}$ and $V^{\prime}\cup\{v\}$ (resp. $V^{\prime}\setminus\{v\}$).
  • By $h(x)$ we denote a witness of $c.e.H.h.c.(d)$ for $\Gamma$. At the end of each step, in Lemma 7.5 we produce $\widehat{h}(x)$ (possibly with some index) which is a witness of $c.e.H.h.c.(d)$ for $\Gamma^{(n)\star}$, defined with the help of $h$.
@@
  • We assume that all of these elements are natural numbers and are ordered according to the standard ordering of the natural numbers.
The main difference with the classical version of the construction from [Dud21] is that instead of $(1,d)$-matchings of infinite graphs we use $(1,d)$-matchings $\mathfrak{M}^{1}_{n}$, $\mathfrak{M}^{2}_{n}$ of finite subgraphs (when we produce $(1,d-1)$-fans to add to the constructed objects). In the Sections 6 and 7 we will show that the choice of the radius allows us to use this finite matchings $\mathfrak{M}^{1}_{n}$, $\mathfrak{M}^{2}_{n}$ without spoiling $c.e.H.h.c.(d)$ in the resulting graphs.
Before describing the construction in detail, we show the single step of the algorithm at the following scheme. We inform the reader that detailed illustrations of all parts of the step will be given at appropriate places.
@@ 5. Construction
From now on in this section we fix a highly computable bipartite graph $\Gamma=(U,V,E)$ which satisfies $c.e.H.h.c.(d)$ and the following conditions:
@@
Let $\Gamma^{\prime}=(U^{\prime},V^{\prime},E^{\prime})$ be a subgraph of $\Gamma=(U,V,E)$ which is also a highly computable bipartite graph satisfying $c.e.H.h.c.(d)$. Then there is an uniform algorithm as follows.
Having an input $u^{*}\in U^{\prime}$ and an odd radius $r$ let $\mathcal{B}(u^{*})$ and $\mathcal{S}(u^{*})$ denote the corresponding ball and the sphere of radius $r$ in $\Gamma$. Put $\mathcal{B}^{\prime}(u^{*})=\mathcal{B}(u^{*})\cap\Gamma^{\prime}$ 111note that $\mathcal{B}^{\prime}(u^{*})$ does not have to be a ball in $\Gamma^{\prime}$ of the center $u^{*}$.
@@
All elements of $V^{\prime}\cap(\mathcal{B}(u^{*})\setminus\mathcal{S}(u^{*}))$ have the same neighbourhoods in $\Gamma^{\prime}$ and $\mathcal{B}^{\prime}(u^{*})$. By Theorem 3.5 there exists a perfect $(1,d)$-matching in $\Gamma^{\prime}$. Thus there exists a matching in $\mathcal{B}^{\prime}(u^{*})$ which satisfies the conditions of a perfect $(1,d)$-matchings for all $u\in U^{\prime}\cap\mathcal{B}(u^{*})$ and $v\in V^{\prime}\cap(\mathcal{B}(u^{*})\setminus\mathcal{S}(u^{*}))$. Since $\Gamma^{\prime}$ is highly computable and $\mathcal{B}^{\prime}(u^{*})$ is finite we find this matching in a computable way.
We now describe an inductive construction which is the heart of the proof of Theorem 4.1. Every step consists of three parts. Some of them are finished by claims stating that certain graphs satisfy Hall’s $d$-harem condition. These statements support further stages of the construction. We start with the detailed description of the first step of the construction.
@@ 5.4. Step 1, part 3
    • $\dot{v}^{\perp}_{0,i-1}:=v_{i}$ for $2<i<d$.
In either case define the fan $M^{1}_{0}$ as the set of edges $(u^{1}_{0},v^{1}_{0,i})$ for $11\leq i\leq d-$ 1.
@@
5.5. Output of the first step
If $\Gamma^{(1)}$ satisfies $c.e.H.h.c.(d)$, set $\Gamma^{(1)\star}:=\Gamma^{(1)}$ and $U^{(1)\perp}=\emptyset,V^{(1)\perp}=\emptyset$. If $\Gamma^{(1)}$ does not satisfy $c.e.H.h.c.(d)$, set $\Gamma^{(1)\star}:=\Gamma^{\prime}_{0}$ and
@@ 5.5. Output of the first step
5.6. Remarks before step $n+1$
Before description of the $n+1$-st step let us discuss the situation after the $n$-th step.
We have constructed graphs $\Gamma^{(n)}$ and $\Gamma^{(n)\star}$, where $\Gamma^{(n)}$ is $U^{(n)}$-reflected and $\Gamma^{(n)\star}$ satisfies $c.e.H.h.c.(d)$. Also note that since $|U^{(n)\perp}\setminus U^{(n-1)\perp}|\leq 1$, there are at most $n$ roots $u^{\perp}_{i}$ of fans in $U^{(n)\perp}$ (see Section 2.4 for the corresponding definition).
@@ 5.7. Step n+1, part 1
Take $u_{n}$, the first element of the set $U^{(n)}$. In order to define $M^{0}_{n}$ we have two possible cases:
@@
    If $u_{n}\not\in U^{(n)\perp}$, then take the ball $\mathcal{B}^{(n)}(u_{n})$ in $\Gamma^{(n)}$ with the center $u_{n}$ and the radius $\max\{4h(d(3n+1))+3,5\}$. Applying Proposition 5.1 compute $\mathfrak{M}^{1}_{n}$, a perfect $(1,d)$-matching in the ball $\mathcal{B}^{(n)\star}(u_{n}):=\mathcal{B}^{(n)}(u_{n})\cap\Gamma^{(n)\star}$. Verify whether there is $j$ with $(u_{v^{0}_{n,j}},v_{u_{n}})\in\Gamma^{(n)\perp}$. By the definition of $\Gamma^{(n)\perp}$ it can happen for at most one $j$ [^2] Note here that it can also happen that $v_{u_{n}}$ is not even in $\Gamma^{(n)}$. If there is such $j$ then it can be equal to $d$ or not. In the first case set $M^{0}_{n}$ to consist of edges $(u_{n},v^{0}_{n,i})\in\mathfrak{M}^{1}_{n}$ for $i\neq d-1$. In all other possibilities we set $M^{0}_{n}$ to consist of edges $(u_{n},v^{0}_{n,i})\in\mathfrak{M}^{1}_{n}$ for $i\leq d-1$.
@@
Note that as a result in the case of existence of $j$ as above the edge $(u_{n},v^{0}_{n,j})$ is included into $M^{0}_{n}$. Although the final part of the algorithm of (2) can be presented easier, see Figure 4, we do not want to miss this point.
@@
The following claim follows from Lemma 7.1 below.
@@
Let $\Gamma^{(n)\star}(-u_{n})$ be $\Gamma^{(n)}(-u_{n})\cap\Gamma^{(n)\star}$. One of the following holds:
@@
  • we can compute $u_{j}^{\perp}\in U^{(n)\perp}$ such that the graph $\Gamma^{(n)\star}(-u_{n},+u_{j}^{\perp})$ satisfies $c.e.H.h.c.(d)$.
5.8. Output of part 1

### v4

@@ 2.2. Reflections
The following notation substantially simplifies the presentation. For every vertex $v\in V$ there exist a vertex from $U$ which is a copy of $v$ (i.e. the same natural number), we denote it by $u_{v}$. If a vertex $u\in U$ has the copy in $V$ then we denote this copy by $v_{u}$.
@@
The graph $\Gamma=(U,V,E)$ is called $U$-reflected if $V$ is a subset of the right copy of $U$ and for every edge $(u,v)\in E$ with $v_{u}\in V$ the edge $(u_{v},v_{u})$ is also in $E$. If additionally $V$ is the right copy of $U$, then $\Gamma$ is called a fully reflected bipartite graph.
It is worth noting that when $\Gamma^{\prime}=(U^{\prime},V^{\prime})$ is an induced subgraph of an $U$-reflected $\Gamma=(U,V,E)$ such that $V^{\prime}$ is a subset of the right copy of $U^{\prime}$ then $\Gamma^{\prime}$ is $U^{\prime}$-reflected.
Theorem 2.6 below states that in the case of fully reflected bipartite graphs Hall’s harem condition allows us to force some additional properties at the expense of obtaining a perfect (1,(d1))(1,(d-1))-matching instead of a $(1,d)$-matching. This is an intermediate statement that avoids computability issues as much as it is possible, the proof of this theorem is available in [HallControlled]. We will now give necessary details.
@@
Let $f$ be a function. If for some $i\neq 0$ we have $f^{i}(u)=u$ then we will say that $u$ is a periodic pointof $f$. For such $u$ and the smallest $i\neq 0$ with $f^{i}(u)=u$, we say that $\{u,f(u),\ldots,f^{i-1}(u)\}$ is a cycle of $f$. Any (1,(d1))(1,(d-1)) matching can be considered as a (d1)(d-1) to $1$ function $f:\mathbb{N}\rightarrow\mathbb{N}$. Moreover, if this matching is perfect, such a function $f$ is total and surjective. We roughly want to show that given a fully reflected bipartite graph satisfying Hall’s $d$-harem condition, there is a perfect (1,(d1))(1,(d-1)) matching $f:\mathbb{N}\rightarrow\mathbb{N}$, such that for each $u$ there exist $i\geq 0$ such that $f^{i}(u)$ is a periodic point.
@@
Let $f:\mathbb{N}\rightarrow\mathbb{N}$ be a (d1)(d-1) to $1$ function. We say that $f$ has controlled sizes of its cycles if each of the following conditions holds:
@@
The following theorem is proved in [HallControlled].
@@
Then there exist a perfect (1,d1)(1,d-1)-matching of $\Gamma$, which realizes a (d1)(d-1) to $1$ function $f:\mathbb{N}\rightarrow\mathbb{N}$ with controlled sizes of its cycles.
This is the theorem which effective version is the goal of the present paper.
@@
Before we state a computable version of Theorem 2.6 we describe our approach to computability of classical Hall’s $d$-harem theorem from [CPD].
@@
To formulate a computable version of Hall’s Harem theorem we also need the following definition of Kierstead [hak].
@@
Note that a computably bipartite (resp. highly computable) graph $\Gamma$ with infinite $U$ and $V$ can be always presented in the form $(\mathbb{N},\mathbb{N},E)$ with computable $E$. Then induced subgraphs of $\Gamma$ are of the form $(U,V,E\cap(U\times V))$. Assuming that $U$ and $V$ are computable subsets of $\mathbb{N}$, the corresponding subgraph admits a computably bipartite (resp. highly computable) presentation. Typically we would not insist on finding the latter; it would suffice to know that $U$ is taken from the left copy of $\mathbb{N}$, but $V$ is taken from the right one.
Manaster and Rosenstein [MR] showed that there is a computably bipartite, highly computable graph that satisfies Hall’s condition for a bipartite graph to have a $(1,1)$-matching, but has no computable $(1,1)$-matching.
Kierstead [hak] found the natural extension of Hall’s condition, called computable expanding Hall’s condition, which allows us to construct computable $(1,1)$-matchings.
Following [CPD], in this paper we focus on computable perfect $(1,k)$-matchings.
@@
In [CPD] we introduced the following condition, which implies the existence of computable perfect $(1,k)$-matchings in highly computable bipartite graphs. For $k=1$ it was formulated earlier in [hak].
@@
In [CPD] we have proven the following.
@@ 4.1. Main Theorem and the aim of the construction
In the proof of Theorem A we need a computable perfect (1,(d1))(1,(d-1)) matching which realizes a function $f:\mathbb{N}\rightarrow\mathbb{N}$, such that the following conditions are satisfied:
@@
[…]
@@
In fact it is easy to see that a computable (d1)(d-1) to $1$ function with controlled sizes of its cycles (Definition 2.5) satisfies conditions (I)–(III).
@@
Then there exist a computable perfect (1,d1)(1,d-1)-matching of $\Gamma$, which realizes a computable (d1)(d-1) to $1$ function $f:\mathbb{N}\rightarrow\mathbb{N}$ with controlled sizes of its cycles.
The proof of the theorem will be finished in Section 8. The next sections focus on a construction used in it.
@@ 4.2. Notation used in the construction
  • $M_{n-1}$ is a set of (1,d1)(1,d-1)-fans added to $M$ at the end of the $n$-th step. Thus $M=\bigcup\limits_{n=1}^{\infty}M_{n-1}$.
@@
  • $\Gamma^{(n)\perp}$ is a graph with the sets of vertices $(U^{(n)\perp},V^{(n)\perp})$ and the set of edges corresponding to (1,d1)(1,d-1)-fans with roots denoted by $u^{\perp}_{i}\in U^{(n)\perp}$.
@@
  • $\Gamma^{(n)\star}:=(U^{(n)}\setminus U^{(n)\perp},V^{(n)}\setminus V^{(n)\perp})$. We will see that $\Gamma^{(n)\star}$ satisfies Hall’s $d$-harem condition.
@@
  • For any subgraph $\Gamma^{\prime}$ of $\Gamma^{(n)}$ we denote by $\Gamma^{\prime}(-u^{0}_{n},\ldots,-u^{j}_{n})$ the graph obtained from $\Gamma^{\prime}$ by removal of the (1,d1)(1,d-1)-fans of $M_{n}$ with roots $u^{0}_{n},\ldots,u^{j}_{n}$.
  • For any subgraph $\Gamma^{\prime}=(U^{\prime},V^{\prime})$ of $\Gamma^{(n)}$ and any $u_{j}^{\perp}\in U^{(n)\perp}$ we denote by $\Gamma^{\prime}(+u_{j}^{\perp})$ the graph induced in $\Gamma^{(n)}$ by the sets of vertices $U^{\prime}\cup\{u_{j}^{\perp}\}$ and $V^{\prime}\cup\{v^{\perp}_{j,i}:1\leq i\leq d-1\}$.
  • For any subgraph $\Gamma^{\prime}=(U^{\prime},V^{\prime})$ of $\Gamma^{(n)}$ and any vertex $v\in V^{(n)}$ we denote by $\Gamma^{\prime}(+v)$ (resp. $\Gamma^{\prime}(-v)$) the graph induced in $\Gamma^{(n)}$ by the sets of vertices $U^{\prime}$ and $V^{\prime}\cup\{v\}$ (resp. $V^{\prime}\setminus\{v\}$).
  • We denote by $h(x)$ a witness of $c.e.H.h.c.(d)$ for $\Gamma$. At the end of each step we produce $\widehat{h}(x)$ (possibly with some index) which is a witness of $c.e.H.h.c.(d)$ for $\Gamma^{(n)\star}$, defined with the help of $h$. See Lemma 7.5 for details.
@@
  • We assume that all of these elements (i.e. $u_{n},v^{j}_{n,i},u^{j}_{n},v^{\perp}_{j,i},\ldots$ ) are natural numbers and are ordered according to the standard ordering of the natural numbers.
The main difference with the non-effective version of the construction given in [HallControlled] is that instead of $(1,d)$-matchings $\mathfrak{M}^{1}_{n}$, $\mathfrak{M}^{2}_{n}$ introduced in the list above (when we produce (1,d1)(1,d-1)-fans to add to the constructed objects), the circumstances of [HallControlled] allow us to use perfect $(1,d)$-matchings of infinite graphs. So, the construction in that paper is simpler. In the present paper we need some additional effort (mainly in Sections 6 and 7) in order to show that under an appropriate choice of the radius, the finite matchings $\mathfrak{M}^{1}_{n}$ and $\mathfrak{M}^{2}_{n}$ work without spoiling $c.e.H.h.c.(d)$ in the resulting graphs.
Before describing the construction in detail, we show the single step of the algorithm in the picture attached below, see Figure 1. We inform the reader that detailed illustrations of all parts of the step will be given in the appropriate places.
Figure 1. Single step.
@@ 5. Construction
From now on in this section we fix a highly computable bipartite graph $\Gamma=(U,V,E)$ that satisfies $c.e.H.h.c.(d)$ and the following conditions:
@@
Let $\Gamma^{\prime}=(U^{\prime},V^{\prime},E^{\prime})$ be a subgraph of $\Gamma=(U,V,E)$ which is also a highly computable bipartite graph that satisfies $c.e.H.h.c.(d)$. Then there is a uniform algorithm as follows.
Having an input $u^{*}\in U^{\prime}$ and an odd radius $r$ let $\mathcal{B}(u^{*})$ and $\mathcal{S}(u^{*})$ denote the corresponding ball and the sphere of the radius $r$ in $\Gamma$. Put $\mathcal{B}^{\prime}(u^{*})=\mathcal{B}(u^{*})\cap\Gamma^{\prime}$ 111note that $\mathcal{B}^{\prime}(u^{*})$ does not have to be a ball in $\Gamma^{\prime}$ of the center $u^{*}$.
@@
All elements of $V^{\prime}\cap(\mathcal{B}(u^{*})\setminus\mathcal{S}(u^{*}))$ have the same neighborhoods in $\Gamma^{\prime}$ and $\mathcal{B}^{\prime}(u^{*})$. By Theorem 3.5 there exists a perfect $(1,d)$-matching in $\Gamma^{\prime}$. Thus there exists a matching in $\mathcal{B}^{\prime}(u^{*})$ which satisfies the conditions of a perfect $(1,d)$-matchings for all $u\in U^{\prime}\cap\mathcal{B}(u^{*})$ and $v\in V^{\prime}\cap(\mathcal{B}(u^{*})\setminus\mathcal{S}(u^{*}))$. Since $\Gamma^{\prime}$ is highly computable and $\mathcal{B}^{\prime}(u^{*})$ is finite, we find this matching in a computable way.
We now describe an inductive construction that is the heart of the proof of Theorem 4.1. Every step consists of three parts. Some of them are finished by claims stating that certain graphs satisfy Hall’s $d$-harem condition. These statements support further stages of the construction. We start with a detailed description of the first step of the construction.
@@ 5.4. Step 1, part 3
    • $\dot{v}^{\perp}_{0,i-1}:=v_{i}$ for $1<i\leq d$.
In either case define the fan $M^{1}_{0}$ as the set of edges $(u^{1}_{0},v^{1}_{0,i})$ for $1\leq i\leq d-1$.
@@
5.5. The output of the first step
If $\Gamma^{(1)}$ satisfies $c.e.H.h.c.(d)$, set $\Gamma^{(1)\star}:=\Gamma^{(1)}$, $U^{(1)\perp}=\emptyset$ and $V^{(1)\perp}=\emptyset$. If $\Gamma^{(1)}$ does not satisfy $c.e.H.h.c.(d)$, set $\Gamma^{(1)\star}:=\Gamma^{\prime}_{0}$ and
@@ 5.5. The output of the first step
5.6. The situation before step n+1
We have constructed graphs $\Gamma^{(n)}$ and $\Gamma^{(n)\star}$, where $\Gamma^{(n)}$ is $U^{(n)}$-reflected and $\Gamma^{(n)\star}$ satisfies $c.e.H.h.c.(d)$. Since $|U^{(n)\perp}\setminus U^{(n-1)\perp}|\leq 1$, there are at most $n$ roots $u^{\perp}_{i}$ of fans in $U^{(n)\perp}$ (see Section 2.4 for the corresponding definition).
@@ 5.7. Step n+1, part 1
Take $u_{n}$, the first element of the set $U^{(n)}$. In order to define $M^{0}_{n}$ we have two possible cases.
@@
    If $u_{n}\not\in U^{(n)\perp}$, then take the ball $\mathcal{B}^{(n)}(u_{n})$ in $\Gamma^{(n)}$ with the center $u_{n}$ and the radius $\max\{4h(d(3n+1))+3,5\}$. Applying Proposition 5.1 compute $\mathfrak{M}^{1}_{n}$, a perfect $(1,d)$-matching in the ball $\mathcal{B}^{(n)\star}(u_{n}):=\mathcal{B}^{(n)}(u_{n})\cap\Gamma^{(n)\star}$. Let $v^{0}_{n,1},\ldots,v^{0}_{n,d}$ be elements of $V^{(n)\star}$, ordered by its numbers, such that $(u_{n},v^{0}_{n,i})\in\mathfrak{M}^{1}_{n}$ for all $i\leq d$. Verify whether there is $j$ with $(u_{v^{0}_{n,j}},v_{u_{n}})\in\Gamma^{(n)\perp}$. By the definition of $\Gamma^{(n)\perp}$ it can happen for at most one jj. [^2] Note here that it can also happen that $v_{u_{n}}$ is not even in $\Gamma^{(n)}$. If there is such $j$ then it can be equal to $d$ or not. In the first case, set $M^{0}_{n}$ to consist of edges $(u_{n},v^{0}_{n,i})\in\mathfrak{M}^{1}_{n}$ for $i\neq d-1$. In all other possibilities, we set $M^{0}_{n}$ to consist of edges $(u_{n},v^{0}_{n,i})\in\mathfrak{M}^{1}_{n}$ for $i\leq d-1$.
@@
Note that as a result in the case of existence of $j$ as in (2) the edge $(u_{n},v^{0}_{n,j})$ is included into $M^{0}_{n}$. Although the final part of the algorithm of (2) can be presented easier, i.e. as in Figure 5, we do not want to miss this point.
@@
Let $\Gamma^{(n)\star}(-u_{n})$ be $\Gamma^{(n)}(-u_{n})\cap\Gamma^{(n)\star}$. Then one of the following conditions holds:
@@
  • there is $u_{j}^{\perp}\in U^{(n)\perp}$ such that the graph $\Gamma^{(n)\star}(-u_{n},+u_{j}^{\perp})$ satisfies $c.e.H.h.c.(d)$.
There is an algorithm which verifies these conditions and finds $u^{\perp}_{j}$ in the second one.
In case (1), $\Gamma^{(n)\star}(-u_{n})=\Gamma^{(n)\star}$ and the claim is obvious. In case (2) the claim follows from Lemma 7.1 below.
5.8. The output of part 1